# Effective algorithms for tensor train decomposition via the UTV framework

### v1

@@ Appendix A Appendix
The TT-ULV algorithm in the right-to-left sweep is given in Algorithm 4. If the low-rank approximation $\widehat{\mathbf{U}}\widehat{\mathbf{L}}\widehat{\mathbf{V}}$ in step 4 is of size $\widehat{\mathbf{L}}\in\mathbb{R}^{I_{1}\cdots I_{k-1}\times r_{k-1}}$ instead of $\widehat{\mathbf{L}}\in\mathbb{R}^{r_{k-1}\times r_{k-1}}$, that is, the two terms $(\mathbf{U}_{1}\mathbf{L}_{11}+\mathbf{U}_{2}\mathbf{L}_{21})\mathbf{V}^{\top}_{1}$ in the ULV (5), to make the matrix $\widehat{\mathbf{V}}=\mathbf{V}_{1}$ orthogonal to the residual term $\mathbf{E}=\mathbf{U}_{2}\mathbf{L}_{22}\mathbf{V}^{\top}_{2}$, then the results similar to Theorem III.1 and Corollary III.1 also hold for the TT-ULV algorithm 4. But it leads to an extra matrix multiplication $\mathbf{U}_{2}\mathbf{L}_{21}$ in each iteration.
@@
0:  A $d$-dimensional tensor $\mathcal{A}\in\mathbb{R}^{I_{1}\times\cdots\times I_{d}}$, fixed TT-ranks $\{r_{0},r_{1},\dots,r_{d}\}$ with $r_{0}=r_{d}=1$.
0:  Right-orthogonal TT-cores $\mathcal{G}^{(1)},\dots,\mathcal{G}^{(d)}$ of the approximation $\widehat{\mathcal{A}}$ with TT-ranks $r_{k}$’s.
1:  Temporary matrix: $\mathbf{C}=\mathrm{reshape}(\mathcal{A},\left[I_{1}\cdots I_{d-1},I_{d}\right])$, i.e., the $(d-1)$th unfolding matrix $\mathbf{A}_{d-1}$.
2:  for $k=d$ to $2$ do
@@
4:   Compute the truncated rank-$r_{k-1}$ approximation from URV decomposition $\mathbf{C}=\widehat{\mathbf{U}}\widehat{\mathbf{L}}\widehat{\mathbf{V}}^{\top}+\mathbf{E}$, where $\widehat{\mathbf{U}}\widehat{\mathbf{L}}\widehat{\mathbf{V}}^{\top}$ with $\widehat{\mathbf{L}}\in\mathbb{R}^{r_{k-1}\times r_{k-1}}$ is the leading term and $\mathbf{E}$ denotes the residual part in the ULV decomposition (3).
5:   New core: $\mathcal{G}^{(k)}=\operatorname{reshape}\left(\widehat{\mathbf{V}}^{\top},\left[r_{k-1},I_{k},r_{k}\right]\right)$.
6:   $\mathbf{C}=\widehat{\mathbf{U}}\widehat{\mathbf{L}}$.
@@
8:  $\mathcal{G}^{(1)}=\mathbf{C}$.
The theory of rank-revealing UTV decomposition tells us that the term $\mathbf{U}_{2}\mathbf{L}_{21}$ in (5) can be very minor. Hence, as presented in Algorithm 4, the middle matrix $\widehat{\mathbf{L}}$ of the ULV truncation in step 4 only retains the size $r_{k-1}\times r_{k-1}$ to reduce the computational cost. At the moment, the sharp bound in Corollary III.1 cannot be obtained theoretically for Algorithm 4, but a relatively larger bound can be given by a similar analysis of Theorem III.1 and the triangle inequality of Frobenius norm.
Theorem A.1.
@@
Based on the above results, as for the TT-ULV algorithm in right-to-left sweep for a prescribed accuracy $\varepsilon$, if the ULV truncation is as step 4 in Algorithm 4, then the truncated error of each ULV decomposition should be $\delta=\frac{\varepsilon}{d-1}\|\mathcal{A}\|_{F}>\frac{\varepsilon}{\sqrt{d-1}}\|\mathcal{A}\|_{F}$, which may generate the TT-cores with much higher TT-ranks than the TT-ULV algorithm 2 in the left-to-right sweep. On the other hand, it can also set $\delta=\frac{\varepsilon}{\sqrt{d-1}}\|\mathcal{A}\|_{F}$, but the middle matrix of ULV truncation $\mathbf{L}\in\mathbb{R}^{I_{1}\cdots I_{k-1}\times r_{k-1}}$ should retain a larger size, which leads to more computations. Wherefore, we do not strongly recommend the TT-ULV algorithm for computing right-orthogonal cores by a right-to-left sweep. On the contrary, the TT-URV case is well-suited for computing the right-orthogonal cores in a right-to-left sweep.

### v2

@@ 5 Conclusions
The TT-ULV algorithm in a right-to-left sweep is given in Algorithm 4. Regarding the size of the matrix $\widehat{\mathbf{L}}$ in step 4, we consider two distinct cases. If the low-rank approximation $\widehat{\mathbf{U}}\widehat{\mathbf{L}}\widehat{\mathbf{V}}$ in step 4 is of size $\widehat{\mathbf{L}}\in\mathbb{R}^{I_{1}\cdots I_{k-1}\times r_{k-1}}$ instead of $\widehat{\mathbf{L}}\in\mathbb{R}^{r_{k-1}\times r_{k-1}}$, that is, the two terms $(\mathbf{U}_{1}\mathbf{L}_{11}+\mathbf{U}_{2}\mathbf{L}_{21})\mathbf{V}^{\top}_{1}$ in the ULV decomposition (4) are retained to make the matrix $\widehat{\mathbf{V}}=\mathbf{V}_{1}$ orthogonal to the residual term $\mathbf{E}=\mathbf{U}_{2}\mathbf{L}_{22}\mathbf{V}^{\top}_{2}$, then results similar to Theorem 3.1 and Corollary 3.1 also hold for the TT-ULV algorithm 4 in the right-to-left sweep. But this approach leads to extra matrix multiplication $\mathbf{U}_{2}\mathbf{L}_{21}$ in each iteration.
@@
0: A $d$th-order tensor $\mathcal{A}\in\mathbb{R}^{I_{1}\times\cdots\times I_{d}}$, and fixed TT-ranks $\{r_{0},r_{1},\dots,r_{d}\}$ with $r_{0}=r_{d}=1$.
0: Right-orthogonal $T$-cores $\mathcal{G}^{(1)},\dots,\mathcal{G}^{(d)}$ of the approximation $\widehat{\mathcal{A}}$ with TT-ranks $r_{k}$’s.
1: Temporary matrix: $\mathbf{C}=\mathrm{reshape}(\mathcal{A},\left[I_{1}\cdots I_{d-1},I_{d}\right])$, i.e., the (d1)(d-1)th unfolding matrix $\mathbf{A}_{d-1}$.
2: for $k=d$ to 22 do
@@
4:  Compute the truncated rank-$r_{k-1}$ approximation from URV decomposition $\mathbf{C}=\widehat{\mathbf{U}}\widehat{\mathbf{L}}\widehat{\mathbf{V}}^{\top}+\mathbf{E}$, where $\mathbf{E}$ denotes the residual part in the ULV decomposition (6).
5:  New core: $\mathcal{G}^{(k)}=\operatorname{reshape}\left(\mathbf{V}_{1}^{\top},\left[r_{k-1},I_{k},r_{k}\right]\right)$.
6:  $\mathbf{C}=\mathbf{U}_{1}\mathbf{L}_{11}$.
@@
[…]
The theory of rank-revealing UTV decomposition tells us that the term $\mathbf{U}_{2}\mathbf{L}_{21}$ in Eq. (4) can be very minor. Hence, as presented in Algorithm 4, the middle matrix $\widehat{\mathbf{L}}$ of the ULV truncation in step 4 should only retain the $\mathbf{L}_{11}\in\mathbb{R}^{r_{k-1}\times r_{k-1}}$ to reduce the computational cost. In contrast, the sharp bound in Corollary 3.1 cannot be obtained theoretically, since the matrix $\widehat{\mathbf{V}}=\mathbf{V}_{1}$ used to yield the core tensor is no longer orthogonal to the residual part $\mathbf{E}=\mathbf{U}_{2}\mathbf{L}_{21}\mathbf{V}_{1}^{\top}+\mathbf{U}_{2}\mathbf{L}_{22}\mathbf{V}_{2}^{\top}$. In this case, a larger bound can be given by a similar analysis of Theorem 3.1 and the triangle inequality of the Frobenius norm.
Proposition A.1.
@@
Based on the above results, the TT-ULV algorithm in a right-to-left sweep can also be designed for a prescribed accuracy $\varepsilon$. If the ULV truncation in step 4 is taken as $\mathbf{U}_{1}\mathbf{L}_{11}\mathbf{V}^{\top}_{1}$, then the equal truncated error of each ULV decomposition should be $\varepsilon_{k}=\frac{\varepsilon}{d-1}\|\mathcal{A}\|_{F}>\frac{\varepsilon}{\sqrt{d-1}}\|\mathcal{A}\|_{F}$, which may generate the TT-cores with much higher ranks than TT-ULV algorithm 2 in the left-to-right sweep. On the other hand, it can also set $\varepsilon_{k}=\frac{\varepsilon}{\sqrt{d-1}}\|\mathcal{A}\|_{F}$, but the middle matrix of ULV truncation $\widehat{\mathbf{L}}\in\mathbb{R}^{I_{1}\cdots I_{k-1}\times r_{k-1}}$ should retain a larger size, which leads to more computations and storage. Wherefore, we do not strongly recommend the TT-ULV algorithm for computing right-orthogonal cores using a right-to-left sweep. On the contrary, the TT-URV case is well-suited for computing the right-orthogonal cores in a right-to-left sweep; see Section 3.2 for details.